# Investigations of MWISP Clumps: 13CO Clump Source Catalogs and Physical Properties

### v1

@@ C.1 Distance-independent Parameters
The physical properties of molecular gas based on CO observations have been extensively studied in the literatures (e.g. Rigby et al., 2019; Li et al., 2023; Dong et al., 2023; Zhuang et al., 2024; Zhang et al., 2024; Feng et al., 2024). Following these studies, we assume that 12CO emission is optically thick (Yuan et al., 2023). The excitation temperature ($T_{\mathrm{ex}}$) of 12CO can be derived as (Bourke et al., 1997):
@@
The optical depths of 13CO in each pixel of a clump are given by:
@@
The 13CO column densities in each spatial pixel of a clump can thus be expressed as (Bourke et al., 1997; Pineda et al., 2010):
@@
The velocity dispersion ($\sigma_{v}$) of the clump is defined as the intensity-weighted standard deviation of velocities (Jiang et al., 2023), calculated as:
@@
where $I_{i}$ is the intensity of voxel $i$ in the clump mask, and $v_{i}$ is the velocity coordinate of voxel $i$. The line width is estimated as:
@@ C.2 Distance-dependent Parameters
To convert 13CO column density to H2 column densities, we adopt the following abundance ratios: $\left[\rm{}^{12}C/^{13}C\right]=6.21\it R\rm_{GC}+18.71$ ($R_{\rm GC}$ in kpc; Milam et al., 2005), and $\left[\rm H_{2}/^{12}C\right]=1.1\times 10^{4}$ (Frerking et al., 1982).
The H2 column density (Pineda et al., 2010) is calculated as:
@@
The mass of a clump is estimated by:
@@
Assuming the clump is spherical, the equivalent radius is expressed as:
@@
The volume density, in cm-3, is calculated as:
@@
The mass surface density, in $M_{\odot}~\rm pc^{-2}$, is calculated as:
@@
Notably, $N_{\rm H_{2}}$ depends on $R_{\rm GC}$ through the isotope ratio $\left[\rm{}^{12}C/^{13}C\right]$, and since $R_{\rm GC}$ depends on the heliocentric distance $d$, the surface density $\Sigma$ is indirectly distance-dependent. As $R_{\rm GC}$ increases (which typically occurs with increasing $d$ for objects in similar Galactic directions), the $\left[\rm{}^{12}C/^{13}C\right]$ ratio increases, resulting in higher derived $N_{\rm H_{2}}$ values and consequently higher $\Sigma$.
@@
where $R_{\rm eq}$ is in pc, $\Delta v$ is in $\rm km\,s^{-1}$, and the derived $M_{\rm vir}$ is in $M_{\odot}$. The virial parameter is defined as:
@@
Figure 13: Parameter loss example for the clump located in the upper right corner of Figure 1(d). The panels show how physical properties vary with normalized temperature gradient ($T_{\text{Edge}}/T_{\text{max}}$): (a) Flux decreases by 6.5%; (b) $\sigma_{v}$ decreases by 7.6%; and (c) $R_{\text{eq}}$ decreases by 4.0%. Blue dots represent observational measurements sampled at $T_{\text{Edge}}$ intervals of $\sim$ 0.22 K, corresponding to typical instrumental noise level. Red curves show cubic polynomial fits to the data, while green crosses indicate extrapolated values at the theoretical zero-intensity boundary.
Figure 14: Probability density distributions of parameter loss rates for all MWISP clumps. The panels display: (a) Flux loss with a median value of $12.4\%$, (b) $\sigma_{v}$ loss with a median value of $5.7\%$, and (c) $R_{\rm eq}$ loss with a median value of $6.1\%$. The dashed vertical lines indicate the median values.
@@ D.1 Parameter Recovery through Extrapolation Method
All clump detection algorithms inherently suffer from parameter underestimation biases. While the MWISP survey employs the FacetClumps algorithm, which significantly reduces flux loss (Jiang et al., 2023), systematic underestimation still occurs. Figure 13 demonstrates how flux, velocity dispersion ($\sigma_{v}$), and equivalent radius ($R_{\text{eq}}$) vary as functions of normalized temperature gradient ($T_{\text{Edge}}/T_{\text{max}}$). These relationships usually exhibit non-linear behavior that can be modeled using cubic or lower-order polynomial fits.
@@
where $Par_{\text{obs}}$ represents the observed parameter value at detection threshold, and $Par_{\text{ext}}$ represents the extrapolated value at the zero-intensity boundary.
Through the extrapolation method (Rosolowsky & Leroy, 2006), we can recover the true parameters of each clump. For the example shown in Figure 13, we measure losses of $6.5\%$ for flux, $7.6\%$ for $\sigma_{v}$, and $4.0\%$ for $R_{\text{eq}}$. Figure 14 displays probability density distributions of parameter loss rates for all MWISP sample as described in Section 5. The asymmetric shapes of the distributions, particularly for flux, suggest that the underestimation bias varies among different clumps, likely depending on their density profiles and S/N. The median values indicate that flux measurements (-12.4%) are most severely affected by the detection sensitivity limitations, followed by $R_{\text{eq}}$ (-6.1%), while $\sigma_{v}$ (-5.7%) shows the least bias. This is align with expectations, as flux measurements integrate over the entire clump volume and are thus more responsive to emission from diffuse outer regions that fall below detection thresholds.
Figure 15: Scaling relations between extrapolated physical parameters of MWISP. Panel shows: (a) $M$-$R_{\rm eq}$; (b) $n({\rm H_{2}})$-$R_{\rm eq}$; (c) $\alpha_{\rm vir}$-$R_{\rm eq}$; (d) $\sigma_{v}$-$R_{\rm eq}$; (e) $\sigma_{v}$-$M$; (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$; (g) $N(^{13}\mathrm{CO})$-$R_{\rm eq}$; (h) $\alpha_{\rm vir}$-$N(^{13}\mathrm{CO})$; and (i) $\sigma_{v}$-$N(^{13}\mathrm{CO})$. The distributions are shown for four subsets: MWISP complete clumps (cyan contours), all extrapolated MWISP clumps (colored points with green contours), extrapolated CMD-C (blue contours), and extrapolated CMD-H (red contours). For each relation, the best-fit power-law index and corresponding $P$ are indicated in the legend.
@@ D.2 Robustness of Scaling Relations Against Extrapolation Correction
Figure 15 includes the distributions and scaling relations for both the original and extrapolated parameters across our sample categories. When comparing the complete non-extrapolated sample with its extrapolated counterpart, we observe only a minor shift in parameter space, with no significant changes in the fitting coefficients. Most parameters remain statistically equivalent within their respective error margins. The most substantial percentage change occurs in the $\sigma_{v}$-$R_{\rm eq}$ relation, where the $\Delta\alpha_{0.9}$ ratio is approximately 6.57%. In conjunction with Figure 11, Figure 15 further indicates that the fitting coefficients for extrapolated CMD-C and CMD-H samples maintain consistent trends relative to the complete sample, thereby preserving the differences documented in Section 5.3.2.
This consistency between extrapolated and non-extrapolated scaling relations demonstrates an important property of power-law relationships in molecular clump: they are governed primarily by intrinsic physical processes rather than by limitations in detection sensitivity thresholds. While extrapolation yields more accurate absolute values for individual clump parameters, the proportional relationships between these parameters remain remarkably stable. This robustness validates our conclusions regarding systematic differences across different clump populations and their association with varying levels of star formation activity.
@@ Appendix E Clump Examples
In Figure 16, we present five exemplary clumps: the clump corresponding to the maximum intensity, the clump with the maximum mass, the clump with the minimum mass, the most massive maser-hosting clump, and the least massive maser-hosting clump. For each clump, we display their velocity-integrated intensity maps, optical depth maps, excitation temperature maps, and mass maps. The data blocks shown are slightly larger than the minimum bounding cube of the clumps. They represent the typical morphology of clumps in our catalogs. The parameters of these clumps are presented in the tables in Appendix B.
Figure 16: Examples of MWISP clumps. Row 1: Clump corresponding to the maximum intensity (position: [15.017, -0.667, 19.261 km s-1]). Row 2: Clump with the maximum mass. Row 3: Clump with the minimum mass. Row 4: The most massive maser-hosting clump. Row 5: The least massive maser-hosting clump. Columns from left to right: (a) velocity-integrated intensity maps, (b) optical depth maps, (c) excitation temperature maps, and (d) mass maps. Labels show catalog ID, galactic coordinates, total flux, mean optical depth, mean excitation temperature, and total mass.

### v2

@@ C.1 Distance-independent Parameters
The physical properties of molecular gas based on CO observations have been extensively studied in the literatures (e.g., Rigby et al., 2019; Li et al., 2023; Dong et al., 2023; Zhuang et al., 2024; Zhang et al., 2024; Feng et al., 2024). Following these studies, we assume that 12CO emission is optically thick (Yuan et al., 2023). The excitation temperature ($T_{\mathrm{ex}}$) of 12CO can be derived as (Bourke et al., 1997):
@@
The optical depths of 13CO in each pixel of a clump are given by
@@
The 13CO column densities in each spatial pixel of a clump can thus be expressed as (Bourke et al., 1997; Pineda et al., 2010)
@@
The velocity dispersion ($\sigma_{v}$) of the clump is defined as the intensity-weighted standard deviation of velocities (Jiang et al., 2023), calculated as
@@
where $I_{i}$ is the intensity of voxel $i$ in the clump mask, and $v_{i}$ is the velocity coordinate of voxel $i$. The line width is estimated as
@@ C.2 Distance-dependent Parameters
To convert 13CO column density to H2 column densities, we adopt the following abundance ratios: $\left[\rm{}^{12}C/^{13}C\right]=6.21\it R\rm_{GC}+18.71$ ($R_{\rm GC}$ in kiloparsecs; Milam et al., 2005), and $\left[\rm H_{2}/^{12}C\right]=1.1\times 10^{4}$ (Frerking et al., 1982).
The H2 column density (Pineda et al., 2010) is calculated as
@@
The mass of a clump is estimated by
@@
Assuming the clump is spherical, the equivalent radius is expressed as
@@
The volume density, in cm-3, is calculated as
@@
The mass surface density, in $M_{\odot}~\rm pc^{-2}$, is calculated as
@@
Notably, $N_{\rm H_{2}}$ depends on $R_{\rm GC}$ through the isotope ratio $\left[\rm{}^{12}C/^{13}C\right]$, and since $R_{\rm GC}$ depends on the heliocentric distance $d$, the surface density $\Sigma$ is indirectly distance dependent. As $R_{\rm GC}$ increases (which typically occurs with increasing $d$ for objects in similar Galactic directions), the $\left[\rm{}^{12}C/^{13}C\right]$ ratio increases, resulting in higher derived $N_{\rm H_{2}}$ values and consequently higher $\Sigma$.
@@
where $R_{\rm eq}$ is in parsecs, $\Delta v$ is in kilometers per second, and the derived $M_{\rm vir}$ is in $M_{\odot}$. The virial parameter is defined as
@@ Appendix D Quantification and Correction of Parameter Biases Induced by Sensitivity Effects
Figure D.1: Parameter loss example for the clump located in the upper right corner of Figure 1(d). The panels show how physical properties vary with normalized temperature gradient ($T_{\text{Edge}}/T_{\text{max}}$): (a) flux decreases by 6.5%, (b) $\sigma_{v}$ decreases by 7.6%, and (c) $R_{\text{eq}}$ decreases by 4.0%. Blue dots represent observational measurements sampled at $T_{\text{Edge}}$ intervals of $\sim$ 0.22 K, corresponding to the typical instrumental noise level. Red curves show cubic polynomial fits to the data, while green crosses indicate extrapolated values at the theoretical zero-intensity boundary.
Figure D.2: Probability density distributions of parameter loss rates for all MWISP clumps. The panels display: (a) the flux loss with a median value of $12.4\%$, (b) the $\sigma_{v}$ loss with a median value of $5.7\%$, and (c) the $R_{\rm eq}$ loss with a median value of $6.1\%$. The dashed vertical lines indicate the median values.
@@ D.1 Parameter Recovery through Extrapolation Method
All clump detection algorithms inherently suffer from parameter underestimation biases. While the MWISP survey employs the FacetClumps algorithm, which significantly reduces flux loss (Jiang et al., 2023), systematic underestimation still occurs. Figure D.1 demonstrates how flux, velocity dispersion ($\sigma_{v}$), and equivalent radius ($R_{\text{eq}}$) vary as functions of normalized temperature gradient ($T_{\text{Edge}}/T_{\text{max}}$). These relationships usually exhibit nonlinear behavior that can be modeled using cubic or lower-order polynomial fits.
@@
where $Par_{\text{obs}}$ represents the observed parameter value at the detection threshold, and $Par_{\text{ext}}$ represents the extrapolated value at the zero-intensity boundary.
Through the extrapolation method (Rosolowsky & Leroy, 2006), we can recover the true parameters of each clump. For the example shown in Figure D.1, we measure losses of $6.5\%$ for flux, $7.6\%$ for $\sigma_{v}$, and $4.0\%$ for $R_{\text{eq}}$. Figure D.2 displays the probability density distributions of parameter loss rates for the entire MWISP sample as described in Section 5. The asymmetric shapes of the distributions, particularly for flux, suggest that the underestimation bias varies among different clumps, likely depending on their density profiles and S/N. The median values indicate that flux measurements (-12.4%) are most severely affected by the detection sensitivity limitations, followed by $R_{\text{eq}}$ (-6.1%), while $\sigma_{v}$ (-5.7%) shows the least bias. This is aligned with expectations, as flux measurements integrate over the entire clump volume and are thus more responsive to emission from diffuse outer regions that falls below detection thresholds.
Figure D.3: Scaling relations between extrapolated physical parameters of MWISP. Panels show (a) $M$-$R_{\rm eq}$, (b) $n({\rm H_{2}})$-$R_{\rm eq}$, (c) $\alpha_{\rm vir}$-$R_{\rm eq}$, (d) $\sigma_{v}$-$R_{\rm eq}$, (e) $\sigma_{v}$-$M$, (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$, (g) $N(^{13}\mathrm{CO})$-$R_{\rm eq}$, (h) $\alpha_{\rm vir}$-$N(^{13}\mathrm{CO})$, and (i) $\sigma_{v}$-$N(^{13}\mathrm{CO})$. The distributions are shown for four subsets: MWISP complete clumps (cyan contours), all extrapolated MWISP clumps (colored points with green contours), extrapolated CMD-C (blue contours), and extrapolated CMD-H (red contours). For each relation, the best-fit power-law index and corresponding $P$ are indicated in the legend.
@@ D.2 Robustness of Scaling Relations Against Extrapolation Correction
Figure D.3 includes the distributions and scaling relations for both the original and extrapolated parameters across our sample categories. When comparing the complete non-extrapolated sample with its extrapolated counterpart, we observe only a minor shift in parameter space, with no significant changes in the fitting coefficients. Most parameters remain statistically equivalent within their respective error margins. The most substantial percentage change occurs in the $\sigma_{v}$-$R_{\rm eq}$ relation, where the $\Delta\alpha_{0.9}$ ratio is approximately 6.57%. In conjunction with Figure 11, Figure D.3 further indicates that the fitting coefficients for extrapolated CMD-C and CMD-H samples maintain consistent trends relative to the complete sample, thereby preserving the differences documented in Section 5.3.2.
This consistency between extrapolated and non-extrapolated scaling relations demonstrates an important property of power-law relationships in molecular clumps: they are governed primarily by intrinsic physical processes rather than by limitations in detection sensitivity thresholds. While extrapolation yields more accurate absolute values for individual clump parameters, the proportional relationships between these parameters remain remarkably stable. This robustness validates our conclusions regarding systematic differences across different clump populations and their association with varying levels of star formation activity.
@@ Appendix E Clump Examples
In Figure E.1, we present five exemplary clumps: the clump corresponding to the maximum intensity, the clump with the maximum mass, the clump with the minimum mass, the most massive maser-hosting clump, and the least massive maser-hosting clump. For each clump, we display the velocity-integrated intensity maps, optical depth maps, excitation temperature maps, and mass maps. The data blocks shown are slightly larger than the minimum bounding cube of the clumps. They represent the typical morphology of clumps in our catalogs. The parameters of these clumps are presented in the tables in Appendix B.
Figure E.1: Examples of MWISP clumps. Row 1: clump corresponding to the maximum intensity (position: [15.017, -0.667, 19.261 km s-1]). Row 2: clump with the maximum mass. Row 3: clump with the minimum mass. Row 4: the most massive maser-hosting clump. Row 5: the least massive maser-hosting clump. Columns from left to right: (a) velocity-integrated intensity maps, (b) optical depth maps, (c) excitation temperature maps, and (d) mass maps. Labels show the catalog ID, Galactic coordinates, total flux, mean optical depth, mean excitation temperature, and total mass.